# Generalization Error of the Tilted Empirical Risk

### v1

@@ D.2 Information-theoretic bounds: details for unbounded loss
Given Assumption 4.4 and Assumption 4.3, the following upper bound holds on the expected tilted generalization error for $\gamma<0$, $\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&\leq\begin{cases}\exp(-\gamma R_{T})\sqrt{\frac{M_{2}I(H;S)}{n}}+\frac{|\gamma|}{2}M_{2},\qquad&\frac{I(H;S)}{n}\leq\frac{\gamma^{2}M_{2}}{2}\\ \frac{\exp(-\gamma R_{T})}{|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)}+\frac{|\gamma|}{2}M_{2},\qquad&\frac{I(H;S)}{n}>\frac{\gamma^{2}M_{2}}{2}.\end{cases}\end{split}$ (86)
@@
Using Lemma B.7, we have
@@
Using a similar approach as for the proof of Proposition C.9 and applying Proposition 4.10, we obtain
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\begin{cases}\exp(-\gamma R_{T})\sqrt{\frac{M_{2}I(H;S)}{n}},\qquad&\frac{I(H;S)}{n}\leq\frac{\gamma^{2}M_{2}}{2}\\ \frac{\exp(-\gamma R_{T})}{|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)},\qquad&\frac{I(H;S)}{n}>\frac{\gamma^{2}M_{2}}{2}.\end{cases}\end{split}$ (90)
@@
Theorem 4.14.
Given Assumption 4.3 and Assumption 4.4, the following lower bound holds on the expected tilted generalization error for $\gamma<0$, $\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&\geq\begin{cases}-\exp(-\gamma R_{T})\sqrt{\frac{M_{2}I(H;S)}{n}}+\frac{\gamma}{2}M_{2},\qquad&\mbox{ if }\max\big{(}\frac{2I(H;S)}{\gamma^{2}M_{2}},\frac{\gamma^{2}M_{2}I(H;S)}{\exp(2\gamma R_{T})}\big{)}\leq n\\ \frac{\exp(-\gamma R_{T})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}M_{2}}{2}\Big{)}+\frac{\gamma}{2}M_{2},\qquad&\mbox{ if }\min(\frac{\exp(\gamma R_{T})-\frac{\gamma^{2}M_{2}}{2}}{I(H;S)},\frac{2I(H;S)}{\gamma^{2}M_{2}})>n.\end{cases}\end{split}$
@@
The proof is similar to Theorem 4.12 and using Corollary 4.11. ∎
Appendix E Proof and details of Section 5
Proposition 5.3.
Under Assumption 4.1, the difference of the tilted population risk (9) between $\mu$ and $\tilde{\mu}$ is bounded as follows;
$\begin{split}&\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[\exp(\gamma\ell(h,\tilde{Z}))])-\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[\exp(\gamma\ell(h,\tilde{Z}))])\leq\frac{\mathbb{TV}(\mu,\tilde{\mu})}{|\gamma|\exp(\gamma R_{T}^{u})}.\end{split}$ (91)
@@ Appendix E Proof and details of Section 5
We have that
$\begin{split}&\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[\exp(\gamma\ell(h,\tilde{Z}))])-\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[\exp(\gamma\ell(h,\tilde{Z}))])\\ &\overset{(a)}{=}\int_{\mathcal{Z}}\frac{\exp(\gamma\ell(h,z))}{|\gamma|\mathbb{E}_{\tilde{Z}\sim\mu}[\exp(\gamma\ell(h,\tilde{Z}))]}(\tilde{\mu}-\mu)(\mathrm{d}z)\\ &\overset{(b)}{\leq}\frac{\mathbb{TV}(\mu,\tilde{\mu})}{|\gamma|\exp(\gamma R_{T}^{u})}\end{split}$ (92)
where (a) and (b) follow from the functional derivative and Lemma B.13. ∎
Theorem 5.4.
Given Assumptions 4.1, 4.2, 5.1 and 5.2, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$ for $\gamma<0$, then the following upper bound holds on the tilted generalization error $\begin{split}\mathrm{gen}_{\gamma}(h,\hat{S})&\leq 2\exp(-\gamma R_{T}^{s})\sqrt{\frac{M_{2}^{s}(\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma R_{T}^{s})(\log(2/\delta))}{3n\gamma}-\frac{\gamma}{2}M_{2}^{u}-\frac{\exp(-\gamma R_{T}^{u})\mathbb{TV}(\mu,\tilde{\mu})}{\gamma},\end{split}$ where $\hat{S}$ is the training dataset under the distributional shift.
@@
where $I_{5}$, $I_{6}$ and $I_{7}$ can be bounded using Lemma B.7, Proposition 5.3 and Theorem 4.5, respectively. ∎
Appendix F Proofs and details of Section 6
@@
Proposition F.1 (Gibbs posterior).
@@
In addition, using Lemma B.3 for uniform distribution, we have
@@
From Proposition F.2, we can observe that for $\gamma>0$ the upper bound on the tilted generalization error under the Gibbs posterior is tighter than the upper bound on the generalization error of the Gibbs posterior under linear empirical risk given in Aminian et al. (2021a).
@@
Appendix G Other Bounds
@@ G.1 Rademacher Complexity
To provide an upper bound on the tilted generalization error, first, we apply the uniform bound, Lemma B.8, and Talagrand’s contraction lemma (Talagrand, 1996) in order to derive a high-probability upper bound on the tilted generalization error; we employ the notation (107).
Proposition G.1.
@@
where (a) holds due to the Lipschitzness of $\log(x)$ in a bounded interval, (b) holds due to the uniform bound Lemma B.8, (c) and (d) hold due to Talagrand’s contraction Lemma B.9).
@@
Then, we obtain an upper bound on the generalization error by combining Proposition G.1, Massart’s lemma (Massart, 2000) and Lemma B.3.
Theorem G.2.
Under the same assumptions as in Proposition G.1, assuming a finite hypothesis space, the tilted generalization error satisfies with probability at least $(1-\delta)$ that $\begin{split}\mathrm{gen}_{\gamma}(h,S)\leq&\frac{\max(1,\exp(-2\gamma M))}{8\gamma}(\exp(\gamma M)-1)^{2}+2AM_{\ell^{\prime}}B\frac{\sqrt{2\log(\mathrm{card}(\mathcal{H}))}}{n}\\ &+\frac{3(A(\gamma)-1)}{|\gamma|}\sqrt{\frac{\log(1/\delta)}{2n}},\end{split}$ where $A(\gamma)=\exp(|\gamma|M)$ and $B^{2}=\max_{h\in\mathcal{H}}\Big{(}\sum_{i=1}^{n}h^{2}(z_{i})\Big{)}$.
@@
where $\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})$ can be bounded using Proposition 3.2. The second term can be bounded by using Proposition G.1 and Massart’s lemma (Lemma B.10). ∎
@@ G.2 A Stability Bound
Theorem G.3.
Under Assumption 3.1, the following upper bound holds with probability at least $(1-\delta)$ under distribution $P_{S}$, $\begin{split}&\mathbb{E}_{P_{S}}\big{[}\mathrm{gen}_{\gamma}(h_{s}(S),S)\big{]}\\ &\leq\frac{(1-\exp(\gamma M))^{2}}{8\gamma}\Big{(}1+\exp(-2\gamma M)\Big{)}+\exp(|\gamma|M)\mathbb{E}_{P_{S},\mu}[|\ell(h_{s}(S),\tilde{Z})-\ell(h_{s}(S_{(i)}),\tilde{Z})|].\end{split}$ (110)
@@
Using Lemma B.3, we have
$\begin{split}&\mathbb{E}_{P_{S}}\big{[}\mathrm{R}(h_{s}(S),\mu)-\frac{1}{\gamma}\log(\mathbb{E}_{P_{S},\mu}[\exp(\gamma\ell(h_{s}(S),\tilde{Z}))])\big{]}\\ &\leq\frac{-\exp(-2\gamma M)}{2\gamma}\mathrm{Var}_{P_{S},\mu}(\exp(\gamma\ell(h_{s}(S),\tilde{Z})))\end{split}$
@@
Definition G.4 (Uniform Stability).
@@
Remark G.5 (Uniform Stability).
Suppose that the learning algorithm is $\beta$-uniform stable with respect to a given loss function. Then, using Theorem G.3, we have

### v2

@@ E.2 Information-theoretic Bounds
Given Assumption E.1, the following upper bound holds on the expected tilted generalization error for $\gamma<0$,
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&\leq\begin{cases}\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}-\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{-\exp(-\gamma\kappa_{t})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}-\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}\end{cases}\end{split}$
@@
Using Lemma C.9, we have
@@
Using a similar approach as for the proof of Proposition D.9 and applying Proposition E.2, we obtain
$\begin{split}\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H,S})\leq\begin{cases}\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}},\qquad&\frac{I(H;S)}{n}\leq\frac{\gamma^{2}\kappa_{t}^{2}}{2}\\ \frac{\exp(-\gamma\kappa_{t})}{|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)},\qquad&\frac{I(H;S)}{n}>\frac{\gamma^{2}\kappa_{t}^{2}}{2}.\end{cases}\end{split}$ (91)
@@
Remark E.5.
Assuming $\gamma=O(n^{-1/2})$, the upper bound in Theorem E.4 has the convergence rate $O(n^{-1/2})$. Note that the result in Theorem E.4 holds for unbounded loss functions, provided that the second moment of the loss function exists.
A similar approach to Theorem E.4 can be applied to derive an information-theoretical lower bound on the expected tilted generalization error for $\gamma<0$ by applying Corollary E.3.
Theorem E.6.
Given Assumption E.1, the following lower bound holds on the expected tilted generalization error for $\gamma<0$,
$\begin{split}\overline{\mathrm{gen}}_{\gamma}(H,S)&\geq\begin{cases}-\exp(-\gamma\kappa_{t})\sqrt{\frac{\kappa_{t}^{2}I(H;S)}{n}}+\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\max\big{(}\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}},\frac{\gamma^{2}\kappa_{t}^{2}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}\big{)}\leq n\\ \frac{\exp(-\gamma\kappa_{t})}{\gamma}\Big{(}\frac{I(H;S)}{n}+\frac{\gamma^{2}\kappa_{t}^{2}}{2}\Big{)}+\frac{\gamma}{2}\kappa_{t}^{2},\qquad&\mbox{ if }\min(\frac{4\big{(}4\zeta^{-2}-1\big{)}I(H;S)}{\gamma^{2}\kappa_{t}^{2}},\frac{2I(H;S)}{\gamma^{2}\kappa_{t}^{2}})>n.\end{cases}\end{split}$
@@
The proof is similar to Theorem E.4 and using Corollary E.3. ∎
Combining Theorem E.6 and Theorem E.4, we can derive the upper bound on the absolute value of the expected generalization error.
Appendix F Proof and details of Section 5
Proposition 5.2 (Restated).
Under Assumption 4.1 and Assumption 5.1, the difference of the tilted population risk (9) between $\mu$ and $\tilde{\mu}$ is bounded as follows;
$\begin{split}&\Big{|}\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[\exp(\gamma\ell(h,\tilde{Z}))])-\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[\exp(\gamma\ell(h,\tilde{Z}))])\Big{|}\leq\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{(\kappa_{u}-\kappa_{s})}.\end{split}$ (92)
@@ Appendix F Proof and details of Section 5
We have for a fixed $h\in\mathcal{H}$ that
$\begin{split}&\Big{|}\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\mu}[\exp(\gamma\ell(h,\tilde{Z}))])-\frac{1}{\gamma}\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[\exp(\gamma\ell(h,\tilde{Z}))])\Big{|}\\ &\overset{(a)}{=}\Big{|}\int_{0}^{1}\int_{\mathcal{Z}}\frac{\exp(\gamma\ell(h,z))}{|\gamma|\mathbb{E}_{\tilde{Z}\sim\mu_{\lambda}}[\exp(\gamma\ell(h,\tilde{Z}))]}(\tilde{\mu}-\mu)(\mathrm{d}z)\mathrm{d}\lambda\Big{|}\\ &\overset{(b)}{\leq}\frac{\mathbb{TV}(\mu,\tilde{\mu})\exp(|\gamma|\kappa_{s})}{|\gamma|}\int_{0}^{1}\exp(|\gamma|\lambda(\kappa_{u}-\kappa_{s}))\mathrm{d}\lambda\\ &=\frac{\mathbb{TV}(\mu,\tilde{\mu})}{|\gamma|}\frac{\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})}{|\gamma|(\kappa_{u}-\kappa_{s})},\end{split}$ (93)
where (a) and (b) follow from the functional derivative with $\mu_{\lambda}=\tilde{\mu}+\lambda(\mu-\tilde{\mu})$ and Lemma C.16. The same approach can be applied for lower bound. ∎
Proposition 5.3 (Restated).
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$ for $\gamma<0$, then the following upper bound holds on the tilted generalization error $\begin{split}\mathrm{gen}_{\gamma}(h,\hat{S})&\leq 2\kappa_{s}\exp(-\gamma\kappa_{s})\sqrt{\frac{(\log(2/\delta))}{n}}\\ &\qquad-\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma}-\frac{\gamma}{2}\kappa_{u}^{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ where $\hat{S}$ is the training dataset under the distributional shift.
@@
where $I_{5}$, $I_{6}$ and $I_{7}$ can be bounded using Lemma C.9, Proposition 5.2 and Theorem 4.2, respectively. ∎
Remark F.1.
It is noteworthy that the upper bound in Proposition 5.3 can be infinite for $\gamma\rightarrow-\infty$ and $\gamma=0$ and a fixed n. Consequently, there must exist a $\gamma\in(-\infty,0)$ that minimizes this upper bound. To illustrate this point, consider the case where $n\rightarrow\infty$; here, the first and second terms in the upper bound would vanish. Thus, we are led to the following minimization problem:
$\gamma^{\star}:=\mathop{\mathrm{arg}\min}_{\gamma\in(-\infty,0)}\left[-\frac{\gamma}{2}\kappa_{u}^{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})}\right],$ (94)
Proposition 5.4 (Restated).
Given Assumptions  4.1 and 5.1, for any fixed $h\in\mathcal{H}$ and with probability least $(1-\delta)$, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, such that the following upper bound holds on the tilted generalization error $\begin{split}\mathrm{gen}_{\gamma}(h,\hat{S})&\geq-\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}+\frac{4\exp(-\gamma\kappa_{s})(\log(2/\delta))}{3n\gamma(1-\zeta)}\\ &\qquad-\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ where $\hat{S}$ is the training dataset under the distributional shift.
Proof.
where $I_{5}$, $I_{6}$ and $I_{7}$ can be bounded using Lemma C.9, Proposition 5.2 and Theorem 4.3, respectively. ∎
Theorem 5.5 (Restated).
Under the same assumptions in Proposition 5.4, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, and with probability at least $(1-\delta)$, the absolute value of the tilted generalization error under distributional shift satisfies $\begin{split}\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,\hat{S})|&\leq\frac{2\kappa_{s}\exp(-\gamma\kappa_{s})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\frac{4\exp(-\gamma\kappa_{s})B(\delta)}{3n\gamma(1-\zeta)}\\ &\qquad-\frac{\gamma\kappa_{u}^{2}}{2}+\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\big{)}}{(\kappa_{u}-\kappa_{s})},\end{split}$ (95) where $B(\delta)=\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)$.
Proof.
The result follows from combining the results of Proposition 5.3, Proposition 5.4 and applying the uniform bound. ∎
Appendix G Proofs and details of Section 6
@@
Proposition G.1 (Gibbs posterior).
@@
In addition, using Lemma C.3 for uniform distribution, we have
@@
From Proposition G.2, we can observe that for $\gamma>0$ the upper bound on the tilted generalization error under the Gibbs posterior is tighter than the upper bound on the generalization error of the Gibbs posterior under linear empirical risk given in Aminian et al. (2021a).
@@
Appendix H Other Bounds
@@ H.1 Rademacher Complexity
To provide an upper bound on the tilted generalization error, first, we apply the uniform bound, Lemma C.11, and Talagrand’s contraction lemma (Talagrand, 1996) in order to derive a high-probability upper bound on the tilted generalization error; we employ the notation (110).
Proposition H.1.
@@
where (a) holds due to the Lipschitzness of $\log(x)$ in a bounded interval, (b) holds due to the uniform bound Lemma C.11, (c) and (d) hold due to Talagrand’s contraction Lemma C.12).
@@
Then, we obtain an upper bound on the generalization error by combining Proposition H.1, Massart’s lemma (Massart, 2000) and Lemma C.3.
Theorem H.2.
Under the same assumptions as in Proposition H.1, assuming a finite hypothesis space, the tilted generalization error satisfies with probability at least $(1-\delta)$ that $\begin{split}\mathrm{gen}_{\gamma}(h,S)\leq&\frac{\max(1,\exp(-2\gamma M))}{8\gamma}(\exp(\gamma M)-1)^{2}+2AM_{\ell^{\prime}}B\frac{\sqrt{2\log(\mathrm{card}(\mathcal{H}))}}{n}\\ &+\frac{3(A(\gamma)-1)}{|\gamma|}\sqrt{\frac{\log(1/\delta)}{2n}},\end{split}$ where $A(\gamma)=\exp(|\gamma|M)$ and $B^{2}=\max_{h\in\mathcal{H}}\Big{(}\sum_{i=1}^{n}h^{2}(z_{i})\Big{)}$.
@@
where $\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu^{\otimes n})$ can be bounded using Proposition 3.2. The second term can be bounded by using Proposition H.1 and Massart’s lemma (Lemma C.13). ∎
@@ H.2 A Stability Bound
Theorem H.3.
Under Assumption 3.1, the following upper bound holds with probability at least $(1-\delta)$ under distribution $P_{S}$, $\begin{split}&\mathbb{E}_{P_{S}}\big{[}\mathrm{gen}_{\gamma}(h_{s}(S),S)\big{]}\\ &\leq\frac{(1-\exp(\gamma M))^{2}}{8\gamma}\Big{(}1+\exp(-2\gamma M)\Big{)}+\exp(|\gamma|M)\mathbb{E}_{P_{S},\tilde{Z}}[|\ell(h_{s}(S),\tilde{Z})-\ell(h_{s}(S_{(i)}),\tilde{Z})|].\end{split}$ (113)
@@
Using Lemma C.3, we have
$\begin{split}&\mathbb{E}_{P_{S}}\big{[}\mathrm{R}(h_{s}(S),\mu)-\frac{1}{\gamma}\log(\mathbb{E}_{P_{S},\mu}[\exp(\gamma\ell(h_{s}(S),\tilde{Z}))])\big{]}\\ &\leq\frac{-\exp(-2\gamma M)}{2\gamma}\mathrm{Var}_{P_{S},\mu}(\exp(\gamma\ell(h_{s}(S),\tilde{Z}))),\end{split}$
@@
Definition H.4 (Uniform Stability).
@@
Remark H.5 (Uniform Stability).
Suppose that the learning algorithm is $\beta$-uniform stable with respect to a given loss function. Then, using Theorem H.3, we have